OpenStudy (anonymous):

Write a quadratic equation in factored form to model the given problem. Be sure to write the entire equation. The height of a triangle is 4 more than the base, and the area of the triangle is 6 square units. Find the length of the base. Let x = the length of the base.

OpenStudy (anonymous):

i can walk you through this one.

OpenStudy (anonymous):

what is the equation for the area of a triangle?

OpenStudy (anonymous):

well thats my problem, i'm not sure how to write an equation for a triangle... i know x = 2 though

OpenStudy (anonymous):

How do you know that x = 2?

OpenStudy (anonymous):

Do you know the area of a rectangle?

OpenStudy (anonymous):

because the base is x, the height is x+4 , and the area is 6, so by the basexheight/2 = area

OpenStudy (anonymous):

2 * 6 \2 =6

OpenStudy (anonymous):

Yes, the area of a triangle is base x height /2. That's what I was asking for that you said you didn't know

OpenStudy (anonymous):

So that's the first step. We'll say that your base is b, and your height is h. Therefore the area of your triangle = b*h/2

OpenStudy (anonymous):

But they tell you that the height is 4 more than the base. You write that in math as: h= 4+b

OpenStudy (anonymous):

They also tell us that the area which we will call A is 6. So now let's put that all together

OpenStudy (anonymous):

A=b*h/2 (area of a triangle) A=6 Therefore b*h/2 = 6 But h=4+b therefore let's substitute for h in the equation above b*(4+b)/2 = 6 If we expand it, we get: ( b*4 + b*b)/2 =6 To remove 2 from the denominator, we multiply both sides by 2 2*( b*4 + b*b)/2 =6*2 The 2s on the left side cancel out ( b*4 + b*b) =12 4b + b^2=12 Subtract 12 from both sides 4b + b^2 -12 = 0 Rearrange b^2 + 4b -12 = 0 That's your quadratic equation

OpenStudy (anonymous):

But they ask for you to write it in factored form, so you have to factor b^2 + 4b -12 = 0 You do that by finding the two values that will add together to give you +4b and multiply together to give you -12b^2. those two values are +6b and -2b [Check: +6b + (-2b) = +4b, and +6b*-2b = -12b^2 ] So we factor b^2 + 4b -12 = 0 b^2 -2b + 6b -12 = 0 b(b-2) + 6(b-2) = 0 (b-2)(b+6) = 0 that means (b-2) = 0 or (b+6) = 0 If b-2 = 0, therefore b = 2 if b+6 = 0, therefore b = -6 You get 2 answers, but you can't really have a "negative" length, so we know that b =2 is the right answer. that's how you get the 2.

OpenStudy (anonymous):

Oh by the way, the question asks you to says that the base = x (not b), so everywhere you see b, please replace with x lol